OpenStudy (anonymous):

Question 10 of 25 Multiple Choice: Please select the best answer and click "submit." What is the solution to this inequality? 17 + x 33 A. x 16 B. x 50 C. x 50 D. x 16

OpenStudy (anonymous):

17 + x = 33?

OpenStudy (anonymous):

To solve this you subtract 17 from both sides to isolate the x so: 17 - 17 + x = 33 - 17 then you have: x = 33 - 17 so: x = 16

OpenStudy (anonymous):

So 16 is positive? @shoplovepc

OpenStudy (anonymous):

Yes